The rates of return on Cherry Jalopies, Inc., stock over the last five years were 10 percent, 11 percent,-5 percent, 4 percent, and 8 percent. Over the same period, the returns on Straw Construction Companys stock were 16 percent, 18 percent,-4 percent, 1 percent and 16 percent. Calculate the variances and the standard deviations for Cherry and Straw. (Do not round intermediate calculations. Enter your variance as a decimal rounded to 5 decimal places. Enter your standard deviation as a percent rounded to 2 decimal places.) Cherry Straw Variance Standard deviation
Answer #1

Cherry:

Average return=Total return/Total time period

=(10+11-5+4+8)/5=5.6%

Return (Return-Average Return)^2
10 (10-5.6)^2=19.36
11 (11-5.6)^2=29.16
-5 (-5-5.6)^2=112.36
4 (4-5.6)^2=2.56
8 (8-5.6)^2=5.76
Total=169.2%

Standard deviation=[Total (Return-Average Return)^2/(Time Period-1)]^(1/2)

=[169.2/(5-1)]^(1/2)

=[169.2/4]^(1/2)

=6.50%(Approx)

Variance=Standard deviation^2

=0.00423

Straw:

Average return=Total return/Total time period

=(16+18-4+1+16)/5=9.4%

Return (Return-Average Return)^2
16 (16-9.4)^2=43.56
18 (18-9.4)^2=73.96
-4 (-4-9.4)^2=179.56
1 (1-9.4)^2=70.56
16 (16-9.4)^2=43.56
Total=411.2%

Standard deviation=[Total (Return-Average Return)^2/(Time Period-1)]^(1/2)

=[411.2/(5-1)]^(1/2)

=[411.2/4]^(1/2)

=10.14%(Approx)

Variance=Standard deviation^2

=0.01028

Cherry Straw
Variance 0.00423 0.01028
Standard deviation 6.50% 10.14%(Approx).